donnaydollbaby:

what is the force required to propel a 2.00 kg object at a rate of 2.71 * 10^5 m/s^s

kittybasil:

\[\Huge{F=m\cdot a}\] Force equals mass times acceleration Force = ?? Mass = 2.00 kg Acceleration = 9.8 m/s^2

kittybasil:

Oh wait

kittybasil:

Acceleration is the rate you're given, sorry

kittybasil:

\[\Huge{F=m\cdot a}\] Force equals mass times acceleration Force = ?? Mass = 2.00 kg Acceleration = \(2.71\cdot10^{5}m/s^{2}\)

kittybasil:

The last one is the correct one. Apologies

donnaydollbaby:

1 attachment
donnaydollbaby:

@kittybasil

kittybasil:

?

donnaydollbaby:

do u see the pic help plz

kittybasil:

Sorry but I am in class. That was the best I could do

jhonyy9:

i think in case of second exercise us formula F = m*g - where g=9,6 m/s^2

jhonyy9:

@kittybasil

jhonyy9:

@ThisGirlPretty your opinion please ? ty